OpenStudy (anonymous):

Solve. x-7/2 = - 4 A: X = -22 B: X = -15 C: X = -1 D: X = 6

OpenStudy (anonymous):

What is your guess?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

(x-7)^2=-4 x^2-14x+49=-4 x^2-14x+49+4=0 x^2-14x+53=0 Δ=b^2-4ac Δ=196-212 Δ=-16

OpenStudy (anonymous):

or c i would go wit b

OpenStudy (anonymous):

Thank you @chordae (:

OpenStudy (anonymous):

welcome